OpenStudy (anonymous):

A current of 0.2 A flows through a wire whose ends are at a p.d of 15V Find resistance heat produced in 1 min

OpenStudy (anonymous):

@Michele_Liano

OpenStudy (michele_laino):

we have to apply the Ohm's law, so the requested resistance R is: \[\Large R = \frac{V}{I} = \frac{{15}}{{0.2}} = ...ohms\]

rvc (rvc):

Yes V = IR

OpenStudy (anonymous):

75?

OpenStudy (michele_laino):

correct!

rvc (rvc):

Yep all the best

OpenStudy (michele_laino):

the heat energy E produced, is given by the subsequent formula: \[\Large E = VI\Delta t = 15 \times 0.2 \times 1 \times 60 = ...Joules\]

OpenStudy (anonymous):

180J

OpenStudy (michele_laino):

that's right!

rvc (rvc):

yay

OpenStudy (anonymous):

Thanks for all the help.

OpenStudy (michele_laino):

:) @rvc

OpenStudy (michele_laino):

thanks! @BlackDranzer

OpenStudy (anonymous):

:) @Michele_Laino